OpenStudy (anonymous):

Find the exact value of the real number y. y = csc^-1(1)

OpenStudy (anonymous):

help me!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

@RadEn do you get this one??:0

OpenStudy (anonymous):

@zepdrix do you know this??

OpenStudy (anonymous):

@aaronq do you have any idea??

OpenStudy (anonymous):

@SithsAndGiggles ............wanna help?

OpenStudy (anonymous):

please help:(

OpenStudy (anonymous):

I see you messaged everyone you know.

OpenStudy (anonymous):

haha yes i was desperate :( but no one helped me though....;(

OpenStudy (anonymous):

Just type in arccos(1) into wolframalpha.com

OpenStudy (anonymous):

It's a pretty good website when it comes to math.

OpenStudy (anonymous):

ohhh they said 0 but i need it in a pi form.......@LordVictorius

OpenStudy (anonymous):

@LordVictorius

OpenStudy (anonymous):

It gives it in that form.

OpenStudy (anonymous):

Try it.

zepdrix (zepdrix):

\[\large y=arccsc(1) \qquad \rightarrow \qquad \csc(y)=1\] Recall that \(\large csc(x)=\dfrac{1}{sin(x)}\) Which gives us,\[\large \frac{1}{\sin(y)}=1 \qquad \rightarrow \qquad \sin(y)=1\]At what angle does the sine function give us a value of 1?

zepdrix (zepdrix):

I'm not sure how you came up with zero, did you punch it into Wolfram correctly? :O http://www.wolframalpha.com/input/?i=y%3Darccsc%281%29

OpenStudy (anonymous):

@zepdrix um....pi or 2pi?

zepdrix (zepdrix):

Don't you remember your special angles? D: Go back to your unit circle, you must have one lying around somewhere!! c:

OpenStudy (anonymous):

Type, "Unit Circle" into google.

OpenStudy (anonymous):

@zepdrix ohhhhh pi/2 right?

zepdrix (zepdrix):

Yes c: good.

OpenStudy (anonymous):

@LordVictorius okay:D

OpenStudy (anonymous):

@zepdrix that's the answer?

zepdrix (zepdrix):

Yes.

OpenStudy (anonymous):

thankyouuuuuu!

OpenStudy (anonymous):

np